# Benchmarking End-To-End Performance of AI-Based Chip Placement Algorithms

### v1

@@ 1 Introduction
We summarize our major contributions as follows. (1) Our proposed ChiPBench is a reproducible and unified evaluation framework of existing AI-based chip placement algorithms using end-to-end performance with fully open source EDA tools. This can effectively facilitate research in chip placement within the AI community. (2) We collected $20$ circuits from various domains, and construct a dataset by executing the EDA workflow from the verilog source code, preserving sufficient physical implementation kits for end-to-end performance evaluation. (3) We evaluate six state-of-the-art AI-based chip placement algorithms, including most popular AI-based chip placement algorithms. (4) Experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics.
@@ 2 Related Work
Datasets Some well-known EDA conferences, such as ISPD and ICCAD, host contests addressing EDA challenges and offer benchmarks with processed data for researchers. However, in the early years (e.g., ISPD2005 (Nam et al., 2005) and ICCAD2004 (Adya et al., 2009)), the provided datasets used overly simplified Bookshelf formats, which are abstracted versions of the actual design kits. Therefore, we cannot evaluate the final PPA of the placement results on those datasets. Recently, ISPD2015 (Bustany et al., 2015) and ICCAD2015 (Kim et al., 2015) have offered benchmarks and datasets closer to real-world applications, including necessary netlist, library, and design exchange files, broadening their utility slightly. Nevertheless, they still lack the essential information (e.g., necessary design kits) to run the open-source EDA tools such as OpenROAD (Kahng & Spyrou, 2021). Beyond these conferences, some other datasets have been developed in various directions. For example, the EPFL (Amarú et al., 2015) benchmarks and the larger OpenABC-D Chowdhury et al. (2021) dataset concentrated on synthetic netlists, primarily for testing modern logic optimization tools with a focus on logic synthesis. CircuitNet 2.0 Jiang et al. (2023), on the other hand, shifted the focus towards providing multi-modal data for prediction tasks, enhancing the capability for various prediction tasks through the use of diverse data modalities. Compared with previous efforts, our proposed dataset focuses on the entire EDA workflow. It provides complete files for each case and necessary design kits, such as timing constraints, library files, and LEF files, offering a comprehensive dataset that supports all stages of design and fosters a more integrated approach to chip design and evaluation.
@@ 3 Background on Electronic Design Automation
Electronic Design Automation (EDA) is a suite of software tools vital for designing and developing electronic systems, primarily integrated circuits (ICs). These tools enable electrical engineers to efficiently transform innovative concepts into functional products, addressing the increasing complexity and demands of modern chip design. EDA optimizes the entire design process from schematic capture to layout and fabrication, reducing time-to-market and enhancing design precision and sophistication. In the chip design workflow, EDA tools support various functions: they perform simulations to verify circuit behavior, execute synthesis to convert high-level descriptions to gate-level implementations, and manage physical layouts to ensure designs can be realized in silicon. As shown in Figure 1, the EDA design flow includes several key stages: logic synthesis  floorplanning, placement, Clock Tree Synthesis (CTS), and routing. Below are concise descriptions of each stage, illustrating their importance in the integrated circuit development process.
@@
Logic Synthesis transforms a high-level circuit description into an optimized gate-level netlist implementation (Berndt et al., 2022; Wang et al., 2024a; b). Floorplan involves deciding the layout of major components within an integrated circuit. During this stage, large blocks and core components are positioned in a way that balances signal integrity, power distribution, and area utilization. Placement involves assigning specific locations to various circuit components within the core area of the chip, following the floorplanning stage. The primary objective of this stage is to strategically place the components to optimize performance metrics such as delay and power consumption while ensuring adherence to design rules (Geng et al., 2024). Clock Tree Synthesis(CTS) creates a clock distribution network within an IC to minimize those clock effects, and ensure the correct timing synchronization for circuit operation. Routing involves creating the physical paths for electrical connectivity between various components on the IC as per the netlist. This stage must handle multiple layers of the chip, avoid obstacles, manage signal integrity, and meet all electrical and timing constraints (Cheng et al., 2022).
@@
Chip Placement The placement stage is crucially divided into two distinct phases: macro placement and cell placement. (1) Macro placement is a critical very large-scale integration (VLSI) physical design problem that targets the arrangement of larger components, such as SRAMs and clock generators—often called macros. This phase significantly impacts the chip’s overall floorplan and essential design parameters like wirelength, power, and area. (2) Following this, the standard cell placement phase addresses the arrangement of the more numerous and smaller standard cells, which serve as the fundamental building blocks of digital designs. This phase typically utilizes analytical solvers to secure an optimized configuration that not only minimizes wirelength but also enhances the electrical performance of the chip.
@@ 4.1 Description of Designs
Due to the oversimplification of datasets in early years, there exists a significant gap between these datasets and real-world applications. For instance, the usually used Bookshelf format (Nam et al., 2005; Adya et al., 2009) is overly simplified so that placement results given in such format are inapplicable for the subsequent stages to obtain a valid final design. Some later datasets (Kim et al., 2015) provide the LEF/DEF and necessary files for running these stages, but the contained circuits are still limited and they still lack some information for open-source tools like OpenROAD to work. For instance, the library file lacks buffer definitions, which is necessary for the clock tree synthesis phase, and the LEF file has incomplete layer definitions, which hinders the routing phase. To address this issue, we construct a dataset with comprehensive physical implementation information across the entire flow. Our dataset involves collecting a series of designs spanning various domains, including components such as CPUs, GPUs, network interfaces, image processing technologies, IoT devices, cryptographic units, and microcontrollers. Additionally, the dataset features a diverse array of sizes, with cell ranges from thousands to nearly a million. The statistics for each case is detailed in Table 1, and we defer more details to Table 4 in Appendix C.
@@
1 CAN-Bus (Tommydag, ) 815 935 0 2637 13 935
2 iot shield (brmcfarl, ) 904 1006 0 2995 33 974
3 FPGA-CAN (WangXuan95, ) 140848 178913 0 532024 4 176472
4 subrisc (Hara-Laboratory, ) 859382 1103295 0 3359066 34 1092653
5 sha256 (secworks, ) 10120 12283 0 38758 77 12176
@@ 3 FPGA-CAN (WangXuan95, ) 140848 178913 0 532024 4 176472
10 camera interface (AngeloJacobo, ) 332 340 0 979 29 316
11 image processing (suntodai, ) 2333 3270 0 7823 64 3170
@@
15 bp_fe (The-OpenROAD-Project, ) 33206 36379 11 111510 2511 36203
16 chameleon (The-OpenROAD-Project, ) 61192 3396 4 10539 145 3285
17 riscv32i (The-OpenROAD-Project, ) 12740 10947 6 35370 135 9875
@@
19 toygpu (openrisc, ) 368081 466513 0 1399167 11 461675
20 mor1kx (matt-kimball, ) 104293 130743 0 374983 576 125979
@@ 5.1 Black-Box-Optimization (BBO) Methods
WireMask-EA (Shi et al., 2023) is a BBO framework recently appearing on the AI conference NeurIPS’2023. It leverages wiremask to greedily guide the mapping from genotypes to phenotypes. Here the concept of wiremask was first proposed by Lai et al. (2022), and it is a matrix representing the potential increases in HPWL when placing the next macro on the canvas. Building on this, Shi et al. (2023) implement different kinds of BBO algorithms, including random search (RS), evolutionary algorithm (EA), and Bayesian optimization (BO).
@@ 5.2 Analytical (Gradient-Based) Methods
AutoDMP (Agnesina et al., 2023) extends DREAMPlace by automating hyperparameter tuning through multi-objective Bayesian optimization. It further speeds up the optimization process and reduces the heavy manual tuning efforts. At that time, this work showcased the promising potential of integrating GPU-accelerated algorithms with machine learning techniques for the automation of VLSI design.
@@ 6.3 Experimental Setup
We apply the aforementioned workflow to evaluate six macro placement algorithms: SA, WireMask-EA, MaskPlace, ChiPFormer, and the default algorithm in OpenROAD. As most of these methods only support the circuit data in a BookShelf format, while the circuits in our used dataset are in a standard LEF/DEF, we start by converting the LEF/DEF files from the floorplan stage of our dataset to BookShelf format to serve as the input for the placement algorithms. After finishing the macro placement stage, we use DREAMPlace to finish the placement of the standard cells. In this stage it conduct two different operations, coarse placement and legalization, with the coarse placement splitting the cells simultaneously using gradient descent algorithms and the legalization stage correct the positions of cells subject to the non-overlapping constraint. The resulting placement files in BookShelf format are then converted back to DEF format using Dreamplace and reintroduced into the original flow. Finally, we report the final metrics, obtaining end-to-end evaluation results. Additionally, we perform placement using OpenRoad’s native Place method and complete the entire flow to obtain the final metrics for comparison with other algorithms.
@@ 7.1 Main Results
We evaluate the AI-based chip placement algorithms, including SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer, using both intermediate metrics and end-to-end performance. The results for macro placement are in Table 2. ChiPFormer and WireMask-EA demonstrated a significant reduction in MacroHPWL compared to OpenROAD using TritonMacroPlacer. WireMask-EA achieved the best performance in terms of MacroHPWL. While these AI-based placement algorithms showed good performance on several intermediate metrics, they perform poorly in terms of the end-to-end metrics compared to OpenRoad, particularly in Power, TNS, and Area. This outcome revealed a significant gap between the originally optimized MacroHPWL intermediate metrics and the final design PPA.
Table 2: The evaluation results of AI-based macro placement algorithms. MacroHPWL, HPWL, Wirelength are intermediate metrics, and the other metrics evaluate the end-to-end performance.
@@
WireMask-EA 0.698 1.115 1.165 1.170 1.058 1.092 1.096 0.865 1.007
SA 0.938 1.028 1.089 1.087 1.045 1.189 1.799 1.136 1.009
DREAMPlace 0.946 1.092 1.161 1.176 1.045 0.988 1.368 1.043 1.006
AutoDMP 0.762 0.950 0.981 0.981 1.021 0.986 1.089 0.973 1.005
MaskPlace 1.744 1.153 1.146 1.149 1.053 1.185 1.462 1.044 1.009
ChiPFormer 0.728 1.078 1.118 1.117 1.061 1.036 2.097 1.222 1.010
@@
Cell Placement As shown in Table 3, DREAMPlace achieved the best results in the intermediate metrics of HPWL, and performed well in terms of Power. However, OpenROAD achieved the best results in WNS, TNS, NVP, and area, further demonstrating the inconsistency between intermediate metrics and final PPA results.
@@ 7.2 Correlation Analysis
In addition to Wirelength, the final PPA metrics of the chip are associated with WNS/TNS, congestion, area, and power. The analyses shows that the correlation between MacroHPWL and these metrics is weak, indicating that optimization of MacroHPWL has minimal impact on these performance indicators. Moreover, the results in Figure 4 show that Wirelength exhibits weak correlations with WNS and TNS as well. This implies that even if a single-point algorithm successfully optimizes metrics such as Wirelength, the ultimate physical implementation might only enhance one aspect of the PPA metrics and may not effectively optimize the other dimensions. Therefore, more appropriate intermediate metrics are needed to better correlate with the actual PPA objectives.

### v2

@@ 1 Introduction
We summarize our major contributions as follows. (1) Our proposed ChiPBench is a reproducible and unified evaluation framework of existing AI-based chip placement algorithms using end-to-end performance with fully open source EDA tools. This can effectively facilitate research in chip placement within the AI community. (2) We collected $20$ circuits from various domains, and construct a dataset by executing the EDA workflow from the verilog source code, preserving sufficient physical implementation kits for end-to-end performance evaluation. (3) We evaluate six state-of-the-art AI-based chip placement algorithms, including most popular AI-based chip placement algorithms. (4) Experiments demonstrate that intermediate metrics have weak correlation with the final design PPA, emphasizing the critical importance of developing algorithms towards optimizing the final design PPA rather than intermediate metrics.
@@ 2 Related Work
Datasets Some well-known EDA conferences, such as ISPD and ICCAD, host contests addressing EDA challenges and offer benchmarks with processed data for researchers. However, in the early years (e.g., ISPD2005 (Nam et al., 2005) and ICCAD2004 (Adya et al., 2009)), the provided datasets used overly simplified Bookshelf formats, which are abstracted versions of the actual design kits. Therefore, we cannot evaluate the final PPA of the placement results on those datasets. Recently, ISPD2015 (Bustany et al., 2015) and ICCAD2015 (Kim et al., 2015) have offered benchmarks and datasets closer to real-world applications, including necessary netlist, library, and design exchange files, broadening their utility slightly. Nevertheless, they still lack the essential information (e.g., necessary design kits) to run the open-source EDA tools such as OpenROAD (Kahng & Spyrou, 2021). Beyond these conferences, some other datasets have been developed in various directions. For example, the EPFL (Amarú et al., 2015) benchmarks and the larger OpenABC-D Chowdhury et al. (2021) dataset concentrated on synthetic netlists, primarily for testing modern logic optimization tools with a focus on logic synthesis. CircuitNet 2.0 (Jiang et al., 2023), on the other hand, shifted the focus towards providing multi-modal data for prediction tasks, enhancing the capability for various prediction tasks through the use of diverse data modalities. Compared with previous efforts, our proposed dataset focuses on the entire EDA workflow. As shown in Table 1, a comparison between our dataset and mainstream datasets highlights the distinctions. It provides complete files for each case and necessary design kits, such as timing constraints, library files, and LEF files, offering a comprehensive dataset that supports all stages of physical implementation and fosters a more integrated approach to chip design and evaluation.
Table 1: Comparison Between Our Dataset and Existing Datasets
Dataset
Complete
Design Suite
Logic Synthesis
Support
Physical Design
Support
Full EDA Flow
Support
Large Scale
&Diversity
ISPD2005
ICCAD2015
EPFL Benchmarks
OpenABC-D
CircuitNet 2.0
ChiPBench (Ours)
@@ 3 Background on Electronic Design Automation
Electronic Design Automation (EDA) is a suite of software tools vital for designing and developing electronic systems, primarily integrated circuits (ICs). These tools enable electrical engineers to efficiently transform innovative concepts into functional products, addressing the increasing complexity and demands of modern chip design. EDA optimizes the entire design process from schematic capture to layout and fabrication, reducing time-to-market and enhancing design precision and sophistication. In the chip design workflow, EDA tools support various functions: they perform simulations to verify circuit behavior, execute synthesis to convert high-level descriptions to gate-level implementations, and manage physical layouts to ensure designs can be realized in silicon. As shown in Figure 1, the EDA design flow includes several key stages: logic synthesis, floorplanning, placement, Clock Tree Synthesis (CTS), and routing. Below are concise descriptions of each stage, illustrating their importance in the integrated circuit development process.
@@
Logic Synthesis transforms a high-level circuit description into an optimized gate-level netlist (Berndt et al., 2022; Wang et al., 2024a; b). Floorplan involves deciding the layout of major components within an integrated circuit, positioning blocks and core components to balance signal integrity, power distribution, and area utilization. Placement involves assigning specific locations to various circuit components within the core area of the chip, following the floorplanning stage. The primary objective of this stage is to strategically place the components to optimize performance metrics such as delay and power consumption while ensuring adherence to design rules (Geng et al., 2024). Clock Tree Synthesis (CTS) creates a clock distribution network within an IC to minimize those clock effects, and ensure the correct timing synchronization for circuit operation. Routing involves creating the physical paths for electrical connectivity between various components on the IC as per the netlist. This stage must handle multiple layers of the chip, avoid obstacles, manage signal integrity, and meet all electrical and timing constraints (Cheng et al., 2022).
@@
Chip Placement The placement stage is crucially divided into two distinct phases: macro placement and cell placement. (1) Macro placement is a critical very large-scale integration (VLSI) physical design problem that targets the arrangement of larger components, such as SRAMs and clock generators—often called macros. This phase significantly impacts the chip’s overall floorplan and essential design parameters like wirelength, power, and area. (2) Following this, the standard cell placement phase addresses the arrangement of the more numerous and smaller standard cells, which serve as the fundamental building blocks of digital designs. This phase typically utilizes analytical solvers to secure an optimized configuration that not only minimizes wirelength but also enhances the electrical and timing performance of the chip.
@@ 4.1 Description of Designs
Due to the oversimplification of datasets in early years, there exists a significant gap between these datasets and real-world applications. For instance, the usually used Bookshelf format (Nam et al., 2005; Adya et al., 2009) is overly simplified so that placement results given in such format are inapplicable for the subsequent stages to obtain a valid final design. Some later datasets (Kim et al., 2015) provide the LEF/DEF and necessary files for running these stages, but the contained circuits are still limited and they still lack some information for open-source tools like OpenROAD to work. For instance, the library file lacks buffer definitions, which is necessary for the clock tree synthesis phase, and the LEF file has incomplete layer definitions, which hinders the routing phase. To address this issue, we construct a dataset with comprehensive physical implementation information across the entire flow. Our dataset involves collecting a series of designs spanning various domains, including components such as CPUs, GPUs, network interfaces, image processing technologies, IoT devices, cryptographic units, and microcontrollers. Additionally, the dataset features a diverse array of sizes, with cell ranges from thousands to nearly a million. Our dataset features a complete design suite that supports the full EDA flow and includes a diverse range of sizes and domains, as illustrated in Table 1. The statistics for each case is detailed in Table 2, and we defer more details to Table 7 in Appendix C.2.
@@
6 bp_fe (The-OpenROAD-Project, ) 33206 36379 11 111510 2511 36203
7 CAN-Bus (Tommydag, ) 815 935 0 2637 13 935
8 DE2_CCD_edge (suntodai, ) 2333 3270 0 7823 64 3170
9 dft48 Brendon Chetwynd, Kevin Bush, Kyle Ingols 48488 52575 68 125501 132 50654
10 FPGA-CAN (WangXuan95, ) 140848 178913 0 532024 4 176472
11 iot shield (brmcfarl, ) 904 1006 0 2995 33 974
12 mor1kx (openrisc, b) 104293 130743 0 374983 576 125979
13 or1200 openrisc (a) 43386 32195 20 97047 383 31958
14 OV7670_i2c (AngeloJacobo, ) 332 340 0 979 29 316
@@
17 sha256 (secworks, ) 10120 12283 0 38758 77 12176
18 subrisc (Hara-Laboratory, ) 859382 1103295 0 3359066 34 1092653
@@
20 toygpu (matt-kimball, ) 368081 466513 0 1399167 11 461675
@@ 5.1 Black-Box-Optimization (BBO) Methods
WireMask-EA (Shi et al., 2023) is a BBO framework that was recently introduced at the NeurIPS 2023 conference, positioning itself as an innovative approach in the intersection of AI and EDA. The framework utilizes a novel concept called wiremask, which plays a crucial role in guiding the mapping process from genotypes to phenotypes in a greedy manner. The wiremask concept was originally introduced by Lai et al. (2022), where it is defined as a matrix that predicts the potential increase in Half-Perimeter Wirelength (HPWL) for each subsequent macro placement on the design canvas. By estimating the wirelength increase, the wiremask helps in making informed decisions during the placement process, thereby potentially improving the quality of the layout. Building upon this concept, Shi et al. (2023) extended the framework to integrate several types of Black-Box Optimization (BBO) algorithms, including random search (RS), evolutionary algorithm (EA), and Bayesian optimization (BO), demonstrating the versatility of the approach in handling complex optimization tasks in chip design.
@@ 5.2 Analytical (Gradient-Based) Methods
AutoDMP (Agnesina et al., 2023) extends DREAMPlace by automating hyperparameter tuning through multi-objective Bayesian optimization. It further accelerates the optimization process and reduces manual tuning efforts. At that time, this work showcased the promising potential of integrating GPU-accelerated algorithms with machine learning techniques for automating VLSI design.
@@ 6.3 Experimental Setup
We apply the aforementioned workflow to evaluate six macro placement algorithms: SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, ChiPFormer, and the default algorithm in OpenROAD. Additionally, we also assess cell placement algorithms for DREAMPlace and AutoDMP. As most of these methods only support the circuit data in a BookShelf format, while the circuits in our used dataset are in a standard LEF/DEF, we start by converting the LEF/DEF files from the floorplan stage of our dataset to BookShelf format to serve as the input for the placement algorithms. After finishing the macro placement stage, the resulting placement files are then converted back to DEF format and reintroduced into the original flow. The resulting placement files in BookShelf format are then converted back to DEF and reintroduced into the original flow. Finally, we report the final metrics, obtaining end-to-end evaluation results. Additionally, we perform global placement and detailed placement using OpenROAD’s native Place method and complete the entire flow to obtain the final metrics for comparison with other algorithms. Our project is open-sourced on GitHub.
@@ 7.1 Main Results
Macro Placement We evaluate the AI-based chip placement algorithms, including SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer, using both intermediate metrics and end-to-end performance. The results for macro placement are in Table 3. ChiPFormer and WireMask-EA demonstrated a significant reduction in MacroHPWL compared to OpenROAD using TritonMacroPlacer. WireMask-EA achieved the best performance in terms of MacroHPWL. While these AI-based placement algorithms showed good performance on several intermediate metrics, they perform poorly in terms of the end-to-end metrics compared to OpenROAD, particularly in Power, TNS, and Area. This outcome revealed a significant gap between the originally optimized MacroHPWL intermediate metrics and the final design PPA.
Table 3: The evaluation results of AI-based macro placement algorithms. MacroHPWL, HPWL, Congestion are intermediate metrics, and the other metrics evaluate the end-to-end performance.
@@
WireMask-EA 0.647 1.027 1.105 1.099 1.015 1.085 0.995 0.967 1.004
SA 0.836 1.044 1.099 1.097 1.062 1.121 1.311 1.109 1.013
DREAMPlace 0.857 0.974 1.049 1.059 1.015 1.112 1.025 1.038 0.999
AutoDMP 0.698 0.892 0.950 0.950 1.013 1.196 1.540 1.176 1.002
MaskPlace 1.681 1.119 1.148 1.148 1.051 1.014 0.978 0.903 1.014
ChiPFormer 0.681 0.976 1.027 1.024 1.015 1.031 1.355 1.223 0.981
@@
Cell Placement As shown in Table 4, DREAMPlace achieved the best results in the intermediate metrics of HPWL, and performed well in terms of Power. However, OpenROAD achieved the best results in WNS, TNS, and NVP, further demonstrating the inconsistency between intermediate metrics and final PPA results.
@@ 7.2 Correlation Analysis
Figure 4: Correlation Between Metrics
In addition to Wirelength, the final PPA metrics of the chip are associated with WNS/TNS, Area, and Power. The analyses shows that the correlation between MacroHPWL and these metrics is weak, indicating that optimization of MacroHPWL has minimal impact on these performance indicators. Moreover, the results in Figure 3 show that Wirelength exhibits weak correlations with WNS and TNS as well. This implies that even if a single-point algorithm successfully optimizes metrics such as Wirelength, the ultimate physical implementation might only enhance one aspect of the PPA metrics and may not effectively optimize the other dimensions. Therefore, more appropriate intermediate metrics are needed to better correlate with the actual PPA objectives.
7.3 Further Analysis
In this section, we analyze specific cases to assess the impact of different methods on PPA using ariane133 as the benchmark. Table 5 presents the experimental data, while Figure 5 shows the performance variations on the worst path caused by different placement algorithms. The AutoDMP method reduces wirelength but worsens timing, with smaller area and lower power. This is due to fewer buffers being added during timing repair, leading to degraded timing despite the reduced area and power. Moreover, wirelength and timing are not always proportional—critical paths still experience delays, indicating that global wirelength optimization may miss timing-sensitive paths, especially in worst-case scenarios. For more details, see Appendix C.1.
Table 5: The evaluation results of ariane133 under AI-based macro placement algorithms.
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1124169 5065453.5 0.226 6583143 0.369 -0.417 -329.353 1970 349228
SA 1683330 5187015.8 0.230 6699484 0.365 -0.512 -650.399 2317 347043
MaskPlace 4444289 6253554.3 0.265 7853892 0.373 -0.349 -244.936 1636 357322
ChiPFormer 1253799 5019138.5 0.226 6581086 0.370 -0.553 -860.952 2703 349005
DREAMPlace 1111023 4826654.0 0.214 6348638 0.367 -0.540 -690.266 2307 348180
AutoDMP 828592 4250870.4 0.192 5694373 0.350 -0.982 -1913.660 3310 344091
OpenROAD 2685856 5260791.9 0.235 6825071 0.359 -0.498 -596.718 2274 352706
Figure 5: The images of the worst path for each method in ariane133.